OpenStudy (anonymous):

which of the following functions decreases, going downward from left ro right? y=2*3^x y=2*.03^X Y=-2*.03^X Y=.02*3^X

OpenStudy (anonymous):

y=2*.03^X if x is set to one, the value of the expressions above are:\[\left\{6,\frac{3}{50},-\frac{3}{50},\frac{3}{5}\right\}\]Not quite sure what the problem is, however, there you are.